# A Modular Approach for Multimodal Summarization of TV Shows

### v2

@@ 1 Introduction
As scene breaks are not always given explicitly, we devise an algorithm to identify them from the order of the speaker names (row 1, Figure 1). Additionally, we select the optimal order in which to re-arrange scenes (row 2, Figure 1), as these often appear in non-linear order (e.g., there can be several subplots or flashbacks). Next, we produce summaries in a two-layer process. A vision-processing module Lei et al. (2020); Lin et al. (2022) converts the video to text using visual captioning (row 3, Figure 1), which leverages the strong specialized ability of vision-to-text models and allows us to treat the problem as one of text-to-text summarization. We also summarize each scene independently with a module specialized for dialogue summarization (row 4, Figure 1). Finally, we use a high-level summarization module specialized for narrative summarization to fuse the sequence of scene summaries into a final summary (row 5, Figure 1).
@@ 3.2 Scene Detection
We define a cost for a given partition, by invoking the minimum description length (MDL) principle Grünwald (2007), in which the optimal representation of a piece of data is that which contains the fewest of bits. Thus, the cost of a partition is the number of bits needed to specify it. We assume the total set of $N$ character names for the entire transcript is given. Then, for each scene, we make a scene-specific codebook for the $n$ characters that appear there, which assigns each character a code as an index from $10,\dots,n-$ 1, which all have length $\leq\lceil\log{n}\rceil$Using exactly $\lceil\log{n}\rceil$ guarantees a prefix-free code Grunwald (1998) [^2] Because we do not need to convert to concrete codes, we omit the ceiling operator in our optimization objective.. The number of possible codebooks with $n$ out of $N$ characters is ($N\choose n$. Imposing e.g., lexicographic order, on these, the number of bits to specify one is
@@
where $N$ is the total number of speakers as before, and $n$ and $l$ are, respectively, the number of distinct speakers and the number of transcript lines in the scene. See Appendix D for details and examples.
@@ 5.2 Results
Aside from ROUGE-1 and ROUGE-Lsum, which are roughly the same for ‘w/o reordering’ and ‘w/o scene-detection’, all ablation settings lead to a drop in all metrics. We provide detailed analysis of the accuracy of the scene detection module in Appendix E.1. Unsurprisingly, the largest drop is in ‘w/o transcript’ as most information is in the dialogue — a TV show without sound or subtitles is difficult to follow. Interestingly, however, this setting still gets some n-grams (ROUGE) and facts (PREFS) correct, showing that our model is extracting useful information from the video. This is also clear from the ‘w/o video’ setting, which shows a drop in all metrics. Qualitatively, we observe that the vision-processing module is most useful for identifying locations, e.g., ‘at the hospital’ which is generally not mentioned in the dialogue. For ‘w/o scene-detection’ and ‘w/o reordering’, the difference is moderate for ROUGE. For PREFS, it is more substantial and highly significant when taken over all facts in the dataset.
@@ Appendix A Example Summaries
Table 3 shows the gold summary for an episode from the SummScreen3D testset. Tables 4 and 57 show the summary of our model, and those of comparison models, respectively, for the same episode.
@@ Appendix E Additional Results
E.1 Accuracy of Scene-detection Algorithm
Here, in Table 9, we report an empirical test of the accuracy of our scene-detection algorithm. These figures are produced using an episode which has explicitly marked scene breaks. We assign each transcript line a label based on the scene it is in with respect to these explicit scene breaks, e.g., all lines in the first scene get the label ‘0’. Then we do the same for the scene breaks produced by our algorithm, and compare the two sets of labels using unsupervised label comparison metrics, as commonly used in clustering problems: accuracy (ACC), normalized mutual information (NMI), and adjusted Rand index (ARI), defined as, e.g., in Sheng and Huber (2020). We compare to two baselines. The first, denoted ‘uniform’, divides each episode into $n$ equally-sized scenes, where $n$ is the average number of scenes per episode in the dataset. The second, denoted ‘uniform oracle’, does the same except sets $n$ to the true number of scenes in that episode.
acc nmi ari
ours 0.890 0.881 0.766
uniform oracle 0.759 0.819 0.538
uniform 0.723 0.809 0.489
E.2 BERTScore

### v3

@@ 1 Introduction
As scene breaks are not always given explicitly, we devise an algorithm to identify them from the order of the speaker names (row 1, Figure 1). Additionally, we select the optimal order in which to re-arrange scenes (row 2, Figure 1), as these often appear in non-linear order (e.g., , cuts between different subplots or flashbacks). Next, we produce summaries in a two-layer process. A vision-processing module Lei et al. (2020); Lin et al. (2022) converts the video to text using visual captioning (row 3, Figure 1), which leverages the strong specialized ability of vision-to-text models and allows us to treat the problem as one of text-to-text summarization. We also summarize each scene independently with a module specialized for dialogue summarization (row 4, Figure 1). Finally, we use a high-level summarization module specialized for narrative summarization to fuse the sequence of scene summaries into a final summary (row 5, Figure 1).
@@ 3.2 Scene Detection
We define a cost for a given partition, by invoking the minimum description length (MDL) principle Grünwald (2007), in which the optimal representation of a piece of data is that which contains the fewest of bits. MDL has been shown to have many nice theoretical properties (Grünwald, 2007) and has been used in recent machine learning applications from interpreting hidden representations Voita and Titov (2020) to measuring noise-free information content in images Mahon and Lukasiewicz (2024). We use it here to formalize the desirable properties of a good segmentation.
Thus, the cost of a partition is the number of bits needed to specify it. We assume the total set of $N$ character names for the entire transcript is given. Then, for each scene, we make a scene-specific codebook for the $n$ characters that appear there, which assigns each character a code as an index from $10,\dots,n-$ 1, which all have length $\leq\lceil\log{n}\rceil$. Using exactly $\lceil\log{n}\rceil$ guarantees a prefix-free code Grunwald (1998) [^2] Because we do not need to convert to concrete codes, we omit the ceiling operator in our optimization objective.. The number of possible codebooks with $n$ out of $N$ characters is ($N\choose n$. Imposing e.g., lexicographic order, on these, the number of bits to specify one is
@@
where $N$ is the total number of speakers as before, and $n$ and $l$ are, respectively, the number of distinct speakers and the number of transcript lines in the scene. See Appendix D for details and examples. Intuitively, the two terms in (1) impose opposite pressures on the solution. Term 1, the bitcost of specifying a new code book, imposes a cost on each scene break, and means the algorithm only inserts a scene break if it leads to a sufficient reduction in term two to outweigh this cost. Term 2, the bitcost of using each code book to specify the characters in the corresponding scene, penalizes solutions in which scenes contain lots of different characters, and encourages instead segmentations in which the number of speaking characters in each scene is minimized.
@@ 5.2 Results
Aside from ROUGE-1 and ROUGE-Lsum, which are roughly the same for ‘w/o reordering’ and ‘w/o scene-detection’, all ablation settings lead to a drop in all metrics. We provide detailed analysis of the accuracy of the scene detection module in Appendix 5.3. Unsurprisingly, the largest drop is in ‘w/o transcript’ as most information is in the dialogue — a TV show without sound or subtitles is difficult to follow. Interestingly, however, this setting still gets some n-grams (ROUGE) and facts (PREFS) correct, showing that our model is extracting useful information from the video. This is also clear from the ‘w/o video’ setting, which shows a drop in all metrics. Qualitatively, we observe that the vision-processing module is most useful for identifying locations, e.g., ‘at the hospital’ which is generally not mentioned in the dialogue. For ‘w/o scene-detection’ and ‘w/o reordering’, the difference is moderate for ROUGE. For PREFS, it is more substantial and highly significant when taken over all facts in the dataset.
5.3 Accuracy of Scene-detection Algorithm
Here, in Table 3, we report an empirical test of the accuracy of our scene-detection algorithm. These figures are produced using an episode which has explicitly marked scene breaks. We assign each transcript line a label based on the scene it is in with respect to these explicit scene breaks, e.g., all lines in the first scene get the label ‘0’. Then we do the same for the scene breaks produced by our algorithm, and compare the two sets of labels using unsupervised label comparison metrics, as commonly used in clustering problems: accuracy (ACC), normalized mutual information (NMI), and adjusted Rand index (ARI), defined as, e.g., in Mahon and Lukasiewicz (2023). We compare to two baselines. The first, denoted ‘uniform’, divides each episode into $n$ equally-sized scenes, where $n$ is the average number of scenes per episode in the dataset. The second, denoted ‘uniform oracle’, does the same except sets $n$ to the true number of scenes in that episode.
acc nmi ari
ours 0.890 0.881 0.766
uniform oracle 0.759 0.819 0.538
uniform 0.723 0.809 0.489
@@ Appendix A Example Summaries
Table 4 shows the gold summary for an episode from the SummScreen3D testset. Tables 5 and 68 show the summary of our model, and those of comparison models, respectively, for the same episode.
@@ Appendix E Additional Results
E.1 BERTScore